# Tiny models from tiny data: Textual and null-text inversion for few-shot distillation

### v1

@@ 4 Related Work
The state-of-the-art performance on miniImageNet [20] has increased steadily over the years [35, 36, 37, 38, 39, 40]. Most recent work rely on other techniques than generative distillation, and often use relatively large models. Regarding tiny models, the TRIDENT method [41] produced good results using a model only 4x as large as the Conv4 model (see Section 5.1), but due to their transductive setting, these results are not directly comparable to our work. Recently, the top performers are models using extra training data [8, 42].
@@
Knowledge distillation as a means of improving accuracy of a baseline model or for model compression has been widely studied [11, 12, 13, 43]. One notable example is SimCLRv2 [44], where a self-supervised model is trained on a large set of unlabeled data, specialized on a small labeled set, and distilled to a potentially smaller model. Like most related methods, the distillation requires a representative unlabeled dataset.
@@
There is a growing interest in using generated training data for supervised learning [45, 46]. A recent work [47] showed that generated data improves imagenet classification, but did not address the few-shot case or transfer to smaller models. Other works use generated data for few-shot classification, but often in the feature domain and without combining it with distillation [48, 49, 50]. Some prior works combine generative models with distillation. One work [51] addresses distillation with limited data using a CVAE and MixUp regularization. However, the authors don’t address the case where the teacher is also a few-shot model, and don’t present few-shot classification results on miniImageNet. One work [52] uses GANs [53] to generate distillation data, also without addressing the few-shot case. Another work [54] concludes that contrastive pre-training is preferable to using GANs. They also include a distillation step in their pipeline, but assume access to a large unlabeled representative set, which we don’t have. In DatasetGAN [55] and related works [56], GANs are used to generate pixelwise labeled distillation training images. In contrast to our work, they require manual annotation of GAN-generated images (rather than using inversion techniques) and a large set of unlabeled application-specific images.
@@
The original textual inversion technique [23] is related to a larger body of literature on text-to-image personalization [57, 58, 59, 60, 61]. Null-text inversion [24] is also related to GAN inversion [62]. Most works in these directions have image generation or editing as the intended application, and don’t show results for few-shot classification.
Most related to our approach, a recent line of work [63, 32] use diffusion models for few-shot learning with good results. As for our work, their few-shot pipeline can be derived back to LabelHalluc [31]. However, those methods require a known text description of the novel classes, and require more computational effort due to running a latent space optimization for every image generation.
@@ 5 Experiments
5.1 Models
As student, we used two of the smallest models that are widely used in the few-shot literature. The Conv4 model is a tiny 4-layer CNN with only 113k parameters that was used already in the original ProtoNet paper [64]. The ResNet12 model (8M parameters) is significantly larger, but still one of the smaller models used frequently in prior art [31, 65, 66, 67, 32]. The main choice of teacher was the P>M>F method [8] due to its simplicity and good performance. Appendix B contains more details.
Figure 5: Accuracy over iterations for two choices of teachers and data sources. PMF denotes the P>M>F teacher [8] (DINO + ProtoNet), and LH denotes the LabelHalluc teacher [31] (ResNet12 + logistic classifier). The left plot is a zoomed-in version of the right plot (note the scale on both axes)
5.2 Ablation experiments
To tune selected hyperparameters, perform ablations and selected comparisons to prior work, we first run experiments with the ResNet12 model over a limited set of of 5 fixed episodes from the validation split of miniImageNet. The same episodes were used in all experiments, ensuring comparable results despite the low number of episodes. Initial hyperparameters were based on previous work [31], with minor adjustment described in Appendix B. For the core TINT method, we found that $\alpha=1$ works best on few-shot classification benchmarks, and used that option for all presented results.
The accuracy over training iterations for a few combinations of data sources and teachers is shown in Figure 5. While the original LabelHalluc [31] accuracy (plot g) saturated after around 200 iterations, the accuracy of our full pipeline (plot a) kept improving. We therefore also tried training cycles up to 10k iterations (dropping the learning rate by a factor 0.1 after 5k iterations). Interestingly, including synthetic data but keeping the original LabelHalluc teacher gave no significant accuracy gain (plot f). Conversely, switching to a stronger teacher without adding more representative data (plot d) did lead to increased accuracy, but with a slow onset. When running for only 300 iterations (as in [31]), one may mistakenly believe that this combination has no potential, whereas it will in fact start to outperform LabelHalluc after around 3k iterations. A plausible explanation is that distillation requires significantly more iterations when transferring knowledge between different architectures. The final accuracy of our method after the full 10k iterations is shown in the top part of Table 2. Note that in order to reach the highest accuracy, both a strong teacher and the synthetic data is required.
Table 2 shows ablations where we kept the few-shot pipeline fixed while simplifying the image generation. Most notably, we run one ablation where the generative model was removed altogether and replaced with real data mined from the LAION dataset [6] (original SD training data) using CLIP embeddings of the class names. Interestingly, this setup performed significantly worse than our full pipeline even though real images and known class names were used. This shows that the generative model contributes with substantial value to the pipeline. More details are provided in Appendix B.5.
Table 1: Comparing teacher and synthetic data options (1k or 4k synthetic examples per class).
@@ 5.2 Ablation experiments
Table 2: Results using various simplifications of the synthetic data aspect of our method.
@@
TI + SR + full $t$ range 93.3
TI + SR + limited $t$ range 93.5
@@
We also compared with data generated by a competing method, NAO+SeedSelect [32]. The authors provided source code for their core image generation method, but not for the few-shot experiments, and we were unable to reproduce their results exactly. Instead, we used their image generation method as a drop-in replacement of TINT in our few-shot pipeline. Their previous results were reported using 1k generated examples, so we used the same setting for TINT in the comparison. The results (bottom part of Table 2) show that TINT performs better in this context. Some additional comparisons using the Conv4 backbone are included in Appendix B.7, due to space constraints. Most notably, using TINT, the Conv4 backbone reached 81.7%, while replacing TINT with NAO+SeedSelect [32] failed to produce good results (67.3%). The latter could likely be improved by hyperparameter tuning, but we did not proceed further in that direction. The take-home message is that it is not straight-forward to get good results using this competing method for tiny students.
Finally, we also compared the TINT method with competing methods using the FID measure [68], reaching 2.7 for the full TINT method, compared to 8.0 for plain textual inversion and 13.0 for NAO+SeedSelect. Due to space constraints, the full results table and experiment details are provided in Appendix B.8.
@@ 5.3 Final evaluation
The final evaluation was made using the test splits of 3 popular few-shot datasets: miniImageNet [20], CUB [69] and CIFAR-FS [70]. Our main results for both student models are summarized in Table 3. The use of additional data is indicated in column Extras, where SD indicates Stable Diffusion [15] trained on the LAION dataset [6] and DINO indicates the pre-trained DINO model [9]. Note that DINO was trained on ImageNet, which causes some semantic overlap between pre-training and novel classes. However, the DINO training is fully unsupervised, and this issue was discussed extensively in the original P>M>F paper [8]. We also acknowledge that a direct comparison between methods with different use of extra data is not fair, as the setting without extra data is of course more challenging. Nevertheless, we chose to include methods that do not rely on extra data in the comparison, to show how large gains that are possible by using pre-trained models in the pipeline.
Table 3: Main results on miniImageNet, CUB and CIFAR-FS, 5-way 5-shot, with 95% confidence intervals. *: Uses textual descriptions of classes and external text-based prior knowledge. **: Results reported by [70]. $\dagger$: Requires text descriptions of the novel classes. $\ddagger$: Uses a pre-trained CLIP backbone and was further trained on ImageNet1k [71], Fungi [72], MSCOCO [73] and WikiArt [74].
Method Model Extras miniImageNet CUB CIFAR-FS
ProtoNet [64] Conv4 - 68.2 $\pm$ 0.66 - 72.0 $\pm$ 0.6 **
R2-D2 [70] Conv4 - 65.4 $\pm$ 0.2 - 77.4 $\pm$ 0.2
MetaQDA [75] Conv4 - 72.64 $\pm$ 0.62 - 77.33 $\pm$ 0.73
BL++/IlluAug[76] Conv4 - - 79.74 $\pm$ 0.60 -
MELR [77] Conv4 - 72.27 $\pm$ 0.35 85.01 $\pm$ 0.32 -
FRN+TDM [78] Conv4 - - 88.89 -
KSTNet [67] Conv4 Text* 73.72 $\pm$ 0.63 - -
TINT (ours) Conv4 DINO, SD 80.29 $\pm$ 1.14 89.56 $\pm$ 0.80 85.43 $\pm$ 1.03
RENet [79] RN12 - 82.58 $\pm$ 0.30 91.11 $\pm$ 0.24 86.60 $\pm$ 0.32
MELR [77] RN12 - 83.40 $\pm$ 0.28 85.01 -
FRN+TDM [78] RN12 - - 92.80 -
LabelHalluc [31] RN12 - 86.54 $\pm$ 0.48 - 90.5 $\pm$ 0.6
KSTNet [67] RN12 Text* 82.61 $\pm$ 0.48 - -
@@
TINT (ours) RN12 DINO, SD 93.13 $\pm$ 0.50 94.57 $\pm$ 0.60 93.18 $\pm$ 0.86
P>M>F [8] ViT/B DINO 98.4 97.0 92.2
@@
Prior top Conv4 results are clustered around 72-74% accuracy, while larger networks routinely reach better performance. This may lead practitioners to believe that model size is the main bottleneck for Conv4, and that larger models are required in order to reach higher performance. Our results show that it is indeed possible to push the performance even for the tiny Conv4 model significantly if extra data is allowed. For the ResNet12 backbone, there are more representative methods to compare with. Our method is on par with NAO+SeedSelect [32] on miniImageNet, but slightly worse on CUB and CIFAR-FS. Note however that their method requires a text description of the classes, which ours does not. On the other hand, we require an additional pre-trained model (DINO) that they do not. Also recall using NAO+SeedSelect as a drop-in replacement of TINT in our few-shot pipeline leads to lower accuracy (see Section 5.2). Since we used their code for the image generation, this discrepancy is likely explained by some aspect of their training setup outside of the core image generation, indicating that even better results could be possible by using TINT under similar conditions.

### v2

@@ 4 Related Work
The state-of-the-art performance on miniImageNet (Vinyals et al., 2016) has increased steadily over the years (Qiao et al., 2018; Oreshkin et al., 2018; Ye et al., 2020; Mangla et al., 2020; Chen et al., 2021; Bateni et al., 2022). Most recent work rely on other techniques than generative distillation, and often use relatively large models. Regarding tiny models, the TRIDENT method (Singh & Jamali-Rad, 2022) produced good results using a model only 4x as large as the Conv4 model (see Section 5.2), but due to their transductive setting, these results are not directly comparable to our work. Nowadays, the top performers are models using extra training data (Hu et al., 2022; Fifty et al., 2024).
@@
Knowledge distillation as a means of improving accuracy of a baseline model or for model compression has been widely studied (Hinton et al., 2015; Gou et al., 2021; Tian et al., 2019; Chen et al., 2022). One notable example is SimCLRv2 (Chen et al., 2020), where a self-supervised model is trained on a large set of unlabeled data, specialized on a small labeled set, and distilled to a potentially smaller model. Like most related methods, the distillation requires a representative unlabeled dataset.
@@
There is a growing interest in using generated training data for supervised learning (Hennicke et al., 2024; Sarıyıldız et al., 2023). Azizi et al. (2023) showed that generated data improves imagenet classification, but did not address the few-shot case or transfer to smaller models. Other works use generated data for few-shot classification, but often in the feature domain and without combining it with distillation (Hariharan & Girshick, 2017; Wang et al., 2018; Xu & Le, 2022). Some prior works combine generative models with distillation. Nguyen et al. (2022) addressed distillation with limited data using a CVAE and MixUp regularization, but do not address the case where the teacher is also a few-shot model, and do not present few-shot classification results on miniImageNet. Ding et al. (2023) uses GANs Goodfellow et al. (2014) to generate distillation data, also without addressing the few-shot case. Saha et al. (2022) concludes that contrastive pre-training is preferable to using GANs. They also include a distillation step in their pipeline, but assume access to a large unlabeled representative set. In DatasetGAN (Zhang et al., 2021c) and related works (Tritrong et al., 2021), GANs are used to generate pixelwise labeled distillation training images. In contrast to our work, they require manual annotation of GAN-generated images (rather than using inversion techniques) and a large set of unlabeled application-specific images.
@@
The original textual inversion method (Gal et al., 2022) is related to a larger body of literature on text-to-image personalization (Gal et al., 2023; Tewel et al., 2023; Cohen et al., 2022; Kumari et al., 2023; Ruiz et al., 2023). Null-text inversion (Mokady et al., 2023) is also related to GAN inversion (Xia et al., 2022). Most works in these directions have image generation or editing as the intended application, and don’t show results for few-shot classification.
A recent line of work (Samuel et al., 2024b; a) use diffusion models for long-tail generation, with few-shot learning as one application examples. As for our work, their few-shot pipeline can be derived back to LabelHalluc Jian & Torresani (2022). Their SeedSelect method (Samuel et al., 2024b) works by optimizing a seed (initial latent noise) to produce images better matching a set of input images. Similarity to target images is measured using direct similarity between the VQ-VAE latents and CLIP embeddings of the generated image and the inputs. With norm-aware optimization (NAO), Samuel et al. (2024a) suggest a way to interpoloate and find centroids of seeds, avoiding areas with unusual norms in the seed/noise space. Taken together, they generate novel images resembling the support images by inverting each support example to the noise side of the diffusion model, computing the centroid using NAO, constructing interpolation paths between the centroid and each inverted support example, picking seeds along these paths, and optimizing each seed using SeedSelect. However, those methods require a known text description of the novel classes, and SeedSelect requires backpropagating the loss through the entire diffusion process, which is very demanding in terms of compute and memory.
Finally, Trabucco et al. (2024) studied using diffusion models for data augmentation, but suggested a different method and evaluated on different datasets. They also made an effort to prevent training data to leak into their evaluation. We note that their data-centric leakage prevention consists of hiding the known class prompts to the diffusion model, which we also do. To illustrate that their method works outside of the domain that stable diffusion was trained on, they evaluate on a new dataset, Leafy Spurge. While the core dataset has been made publicly available (Doherty et al., 2024), their particular setup ($250250x$ 250 regions classified by a domain expert) has not. Therefore, a direct comparison on this dataset is not possible. We do compare to their core image generation method in Section 5.3, but used in our overall pipeline.
@@ 5 Experiments
Table 1: FID measure (64 features) in comparison with NAO+SeedSelect (Samuel et al., 2024a) and for selected ablations.
Method FID $\downarrow$
NAO+SS 13.0
Plain TI 8.0
TI + superres + limited t range 4.0
Full TINT 2.7
Table 2: Comparing teacher and synthetic data options (1k or 4k synthetic examples per class) for our method and NAO+SeedSelect (Samuel et al., 2024a).
@@
5.1 Generated data quality
We first study the data quality of examples generated using TINT compared to competing methods and simplified versions of our own method. To be useful for distillation, we would like the data to be both diverse and representative of the true classes. A popular way of evaluating synthetic data in these respects is using the FID measure (Heusel et al., 2017). We note that the original measure involves estimating a $2048\times 2048$ covariance matrix. This is numerically problematic in our case, since there are only 600 examples per class in miniImageNet. Instead, following Samuel et al. (2024a), we use a feature space of only 64 dimensions, using pooled features from an earlier layer of the Inception V3 model. We generated 300 random images per class, compared to 300 randomly drawn real images of the same class, repeated the process for 5 episodes (with 5 classes per episode) and computed the average FID score. The results are shown in Table 2. Even though this is a rather coarse measure, it shows that the TINT method in some respect generates images following the novel class images distribution better than competing methods.
Our main goal is not to generate data that is pleasing to look at, and we therefore opt against using human assessment as quality metric. Diversity and quality of the generated samples is in our case only important if it helps us improve the final classifier accuracy. For example, if the generated images show very little diversity (e.g. just generating the support examples over and over again with some noise added), this would not help the classifier distillation training much, and the few-shot performance would be poor. In that sense, final classifier accuracy is already an implicit measure of generated data diversity and quality.
For a better understanding of the method limitations, we include an example of a failure case in Figure 5. Here, the inversion procedure fails to capture the essence of the roundworm class from imagenet, and generated images are frequently reminiscent of atmospherical phenomena or pyrotechnics rather than of worms. This is understandable based on the two microscopy images present in the support set. Recall however that the negative effects of non-desired variation in the generated images are mitigated by the use of a teacher.
Handling significant domain shift is out of scope for this work. Nevertheless, we include an example of generated aerial images in Appendix C, to give at least a qualitative indication in this direction.
Figure 5: More qualitative examples of support (top) and generated (bottom) examples for two classes from miniImageNet. Left: guitar, right: roundworm (failure example).
Refer to captionRefer to caption
Refer to captionRefer to caption
5.2 Student and teacher models
As student, we used two of the smallest models that are widely used in the few-shot literature. The Conv4 model is a tiny 4-layer CNN with only 113k parameters that was used already in the original ProtoNet (Snell et al., 2017). The ResNet12 model (8M parameters) is significantly larger, but still one of the smaller models used frequently in prior art (Jian & Torresani, 2022; Xie et al., 2022; Luo et al., 2021; Li et al., 2023; Samuel et al., 2024a). The main choice of teacher was the P>M>F method (Hu et al., 2022) due to its simplicity and good performance. Appendix B contains more details.
Figure 6: Accuracy over iterations for two choices of teachers and data sources. PMF denotes the P>M>F teacher by Hu et al. (2022) (DINO + ProtoNet), and LH denotes the LabelHalluc teacher (Jian & Torresani, 2022) using ResNet12 and logistic classifier. The left plot is a zoomed-in version of the right plot (note the scale on both axes)
5.3 Initial experiments
To tune selected hyperparameters, perform ablations and selected comparisons to prior work, we first run experiments with the ResNet12 model over a limited set of of 5 fixed episodes from the validation split of miniImageNet. The same episodes were used in all experiments, ensuring comparable results despite the low number of episodes. Initial hyperparameters were based on LabelHalluc (Jian & Torresani, 2022) with minor adjustment described in Appendix B. For the core TINT method, we found that $\alpha=1$ works best on few-shot classification benchmarks, and used that option for all presented results.
The accuracy over training iterations for a few combinations of data sources and teachers is shown in Figure 6. While the original LabelHalluc accuracy (plot g) saturated after around 200 iterations, the accuracy of our full pipeline (plot a) kept improving. We therefore also tried training cycles up to 10k iterations (dropping the learning rate by a factor 0.1 after 5k iterations). Interestingly, including synthetic data but keeping the original LabelHalluc teacher gave no significant accuracy gain (plot f). Conversely, switching to a stronger teacher without adding more representative data (plot d) did lead to increased accuracy, but with a slow onset. When running for only 300 iterations (as in Jian & Torresani (2022)), one may mistakenly believe that this combination has no potential, whereas it will in fact start to outperform LabelHalluc after around 3k iterations. A plausible explanation is that distillation requires significantly more iterations when transferring knowledge between different architectures. The final accuracy of our method after the full 10k iterations is shown in the top part of Table 2. Note that in order to reach the highest accuracy, both a strong teacher and the synthetic data is required.
Table 4 shows ablations where we kept the few-shot pipeline fixed while simplifying the image generation. This also includes mimicking the DA-Fusion method (Trabucco et al., 2024), using TI and noising each support example to a random time step before denoising. We also run one ablation where the generative model was removed altogether and replaced with real data mined from the LAION dataset (Schuhmann et al., 2022) (original SD training data) by comparing CLIP embeddings of the class names and LAION images. Interestingly, this setup performed significantly worse than our full pipeline even though real images and known class names were used. This shows that the generative model contributes with substantial value to the pipeline. More details are provided in Appendix B.5.
@@ 5.3 Initial experiments
Table 3: Results using various variations of the synthetic data aspect.
@@
SR + TI + full $t$ range 93.3
SR + TI + limited $t$ range 93.5
SR + DA-Fusion 92.8
@@
Table 4: Selected results for the Conv4 backbone on the 5 fixed episodes, miniImageNet validation split compared to the core ideas used in NAO+SeedSelect (Samuel et al., 2024a) and DA-Fusion (Trabucco et al., 2024)
Method Syn teacher Acc
Plain TI (4k) PMF 70.4
SR + DA-Fusion (4k) PMF 77.4
TINT (4k) PMF 81.7
NAO+SS (1k) None 67.3
NAO+SS (1k) PMF 64.0
TINT (1k) PMF 80.9
We also compared with data generated by th competing NAO+SeedSelect method (Samuel et al., 2024a). The authors provided source code for their core image generation method, but not for the few-shot experiments, and we were unable to reproduce their results exactly. Instead, we used their image generation method as a drop-in replacement of TINT in our few-shot pipeline. Their previous results were reported using 1k generated examples, so we used the same setting for TINT in the comparison. The results (bottom part of Table 2) show that TINT performs better in this context. Some additional comparisons using the Conv4 backbone are shown in Table 4, using the same 5 fixed miniImageNet episodes. No additional hyperparameter tuning was applied here, so the results can likely be improved for both methods. However, the results show that TINT beats textual inversion and generation using the DA-Fusion (Trabucco et al., 2024) and NAO+SeedSelect (Samuel et al., 2024a) ideas in this context.
@@ 5.4 Final evaluation
The final evaluation was made using the test splits of 3 popular few-shot datasets: miniImageNet (Vinyals et al., 2016), CUB (Wah et al., 2011) and CIFAR-FS (Bertinetto et al., 2018). Our main results for both student models are summarized in Table 5. The use of additional data is indicated in column Extras, where SD indicates Stable Diffusion (Rombach et al., 2022) trained on the LAION dataset (Schuhmann et al., 2022) and DINO indicates the pre-trained DINO model (Caron et al., 2021). Note that DINO was trained on ImageNet, which causes some semantic overlap between pre-training and novel classes. However, the DINO training is fully unsupervised, and this issue was discussed extensively by Hu et al. (2022). We also acknowledge that a direct comparison between methods with different use of extra data is not fair, as the setting without extra data is of course more challenging. Nevertheless, we chose to include methods that do not rely on extra data in the comparison, to show how large gains that are possible by using pre-trained models in the pipeline.
Table 5: Main results on miniImageNet, CUB and CIFAR-FS, 5-way 5-shot, with 95% confidence intervals. *: Uses textual descriptions of classes and external text-based prior knowledge. **: Results reported by Bertinetto et al. (2018). $\dagger$: Requires text descriptions of the novel classes. $\ddagger$: Uses a pre-trained CLIP backbone and was further trained on ImageNet1k Deng et al. (2009), Fungi Brigit Schroeder (2018), MSCOCO Lin et al. (2014) and WikiArt Saleh & Elgammal (2015).
Method Extras miniImageNet CUB CIFAR-FS
Conv4 backbone
ProtoNet Snell et al. (2017) - 68.2 $\pm$ 0.66 - 72.0 $\pm$ 0.6 **
R2-D2 Bertinetto et al. (2018) - 65.4 $\pm$ 0.2 - 77.4 $\pm$ 0.2
MetaQDA Zhang et al. (2021b) - 72.64 $\pm$ 0.62 - 77.33 $\pm$ 0.73
BL++/IlluAugZhang et al. (2021a) - - 79.74 $\pm$ 0.60 -
MELR Fei et al. (2020) - 72.27 $\pm$ 0.35 85.01 $\pm$ 0.32 -
FRN+TDM Lee et al. (2022) - - 88.89 -
KSTNet Li et al. (2023) Text* 73.72 $\pm$ 0.63 - -
TINT (ours) DINO, SD 80.29 $\pm$ 1.14 89.56 $\pm$ 0.80 85.43 $\pm$ 1.03
ResNet12 backbone
RENet Kang et al. (2021) - 82.58 $\pm$ 0.30 91.11 $\pm$ 0.24 86.60 $\pm$ 0.32
MELR Fei et al. (2020) - 83.40 $\pm$ 0.28 85.01 -
FRN+TDM Lee et al. (2022) - - 92.80 -
LabelHalluc Jian & Torresani (2022) - 86.54 $\pm$ 0.48 - 90.5 $\pm$ 0.6
KSTNet Li et al. (2023) Text* 82.61 $\pm$ 0.48 - -
@@
TINT (ours) DINO, SD 93.13 $\pm$ 0.50 94.57 $\pm$ 0.60 93.18 $\pm$ 0.86
ViT/B backbone
P>M>F Hu et al. (2022) DINO 98.4 97.0 92.2
@@
Prior top Conv4 results are clustered around 72-74% accuracy, while larger networks routinely reach better performance. This may lead practitioners to believe that model size is the main bottleneck for Conv4, and that larger models are required in order to reach higher performance. Our results show that it is indeed possible to push the performance even for the tiny Conv4 model significantly if extra data is allowed. For the ResNet12 backbone, there are more representative methods to compare with. Our method is on par with NAO+SeedSelect (Samuel et al., 2024a) on miniImageNet, but slightly worse on CUB and CIFAR-FS. Note however that their method requires a text description of the classes, which ours does not. On the other hand, we require an additional pre-trained model (DINO) that they do not. Also recall using NAO+SeedSelect as a drop-in replacement of TINT in our few-shot pipeline leads to lower accuracy (see Section 5.3). Since we used their code for the image generation, this discrepancy is likely explained by some aspect of their training setup outside of the core image generation, indicating that even better results could be possible by using TINT under similar conditions.